OpenStudy (anonymous):

partial derivative f(w,z) = w/(w^2+z^2)

OpenStudy (cruffo):

Are you thinking of using quotient rule or product rule for this one?

OpenStudy (anonymous):

how would you use this as a product rule?

OpenStudy (cruffo):

\[\large f(w,z) =\frac{w}{w^2+z^2} = w\left(w^2+z^2\right)^{-1}\]

OpenStudy (cruffo):

Though I don't think it would be much better that way. Just thought I'de get your take on it.

OpenStudy (anonymous):

i think the whole point of this problem is to remind myself of quotient rule lol

OpenStudy (cruffo):

:) If \(f = \dfrac{u}{v}\), then \(f' = \dfrac{u'v - uv'}{v^2}\)

OpenStudy (anonymous):

yeap so then \[1(w^2+z^2)-w(2w)/ (w^2+z^2)^2\]

OpenStudy (anonymous):

?

OpenStudy (cruffo):

Yes. Though it can be simplified by gathering like terms in the numerator.

OpenStudy (anonymous):

yeah i know

OpenStudy (cruffo):

I figured you did :) But better to mention it than not.

OpenStudy (anonymous):

for fz is -2zw/(w^2+z^2)^2?

zepdrix (zepdrix):

Yah looks good for fz c:

OpenStudy (anonymous):

thanks